use the following results from a test for marijuana use, which is provided by a certain drug testing company. Among 143 subjects with positive test results, there are 20 false positive results, among 158 negative results, there are4 false negative results. if one of the test sibjects is randomly selected, find the probability that the subjecttestes negative or did not use marijuana.
Answer #1

Total number of subject tested

= 143 + 158

= 301

Now number of total negative

= Number of negative results - number of false negative + number of false positive

= 158 - 4 + 20

= 174

Probability that the subject tests negative is 174/301 = 0.578(upto three decimal)